# Quantum enhanced stratification of Breast Cancer: exploring quantum expressivity for real omics data

### v1

@@ II.2 Quantum Embedding of classical data
Where $R_{ZZ}$ is a two qubits rotation that makes them entangled and $S$ is the set of qubits to be entangled. Nevertheless, this choice does not guarantee alone an advantage over classical embedding as it is deeply influenced by the data characteristics and tasks [29]. The specifics of the quantum embeddings chosen for our experiments are described in the next ”Experimental set up” section.
@@ II.3 Expressivity regularization
The inherent high-dimensionality of the Hilbert space used for projecting the input data can hinder the model performance. In particular for QK it has been shown that, as a result of highly expressive problem agnostic encoding, the non-diagonal entries distribution often skews around the mean [35], making the points in the Quantum Hilbert space nearly indistinguishable. This phenomenon has a close parallelism to the curse of dimensionality of classical ML. It has been shown [36, 37] that acting on the parametric rotational gate angles can regularise the model expressivity. This strategy is implemented with a tunable hyperparameter called bandwidth, which is a constant within the range $[0,1]$. This parameter effectively scales the input, confining the embeddings to a more restricted region of the quantum state space as in:
@@ III.1 Breast Cancer Multi-omics dataset
III.2 Quantum encoding of data
@@
The choice of a suitable quantum embedding, in case of complex real data, is not a straightforward task [11]. Nevertheless, it is possible to explore, as set out in this work, a sub set of configurations of the enconding scheme, in particular by varying the entanglement map of the unitary ansatz. In this way, a trade-off between the expressivity of the kernel and the task-related performance, can be assessed. We specifically explored three configuration of the IPQ-style encoding scheme, making use of unitary templates with different levels of entanglement of the qubits. The most simple configuration employed is the Z feature map without any entanglement, utilising a sequence of Hadamard and rotational $P$ gates, a single qubit rotation around the $Z$ axis, implemented in our case for a 4-dim classical input $x$ as shown in Fig. 1(a). We then implemented the ZZ feature map with both linear and full entanglement: in the linear entanglement, each qubit is entangled with its neighbouring qubits, while in the full entanglement, each qubit is entangled with every other qubit. These feature map resulted respectively in the embedding shown in Fig.1(b) and Fig. 1(c). Furthermore, we regularised the expressivity of the quantum embeddings tuning the angle range $\beta$. For each feature map, we tested five different scaling of the angle range so that the input data $x\in[0,\beta]$, where $\beta\in\{\frac{\pi}{8},\frac{\pi}{4},\frac{\pi}{2},\pi,2\pi\}$.
@@ III.3 Clustering evaluation
In which $a(i)$ and $b(i)$ are respectively the average distance of the $i$-th point to its own cluster, and the average distance to its second closest cluster. The value of the both the SC and the $s(i)$ ranges in the interval $\{-1,1\}$ where a high value indicates that the object is well matched to its own cluster and weakly matched to neighbouring clusters. In addition, we employed Adjusted Mutual Information (AMI) as a measure of agreement between two sets of labelled data [43]. Given two partitions $U$ and $V$ of the same set of $N$ points, the AMI between them is calculated as:
@@ IV.1 Exact simulation
We analysed the performance of the chosen quantum embeddings in an ideal noiseless setting on the full data set with 1980 samples and compared the results with a classical RBF embedding. In this phase we computed the QKs on a locally simulated quantum computer with 4 qubits, using the qasm simulator backend from Qiskit , which is designed to simulate the behaviour of an ideal quantum circuit. Considering no assumptions were made regarding the partition of our data, we optimised the number of clusters (K) in a range from 2 to 10, where $K=10$ refer to the IntClust subtypes [17]. Specifically, we evaluate the SC, as defined in eq 11, for each feature map (i.e. quantum and RBF) and the quantum kernels have been evaluated for input values within the range $[0,\beta]$, with $\beta\in\{\frac{\pi}{8},\frac{\pi}{4},\frac{\pi}{2},\pi,2\pi\}$. The results of the SC for the exact simulations are shown in Fig. 2. Overall, it can be noticed that ZZ-linear map outperforms the classical RBF model for $K=2$, the trivial case, and for finer grouping with $K\geq 9$, with $\beta=\frac{\pi}{8}$ and with $\beta=\frac{\pi}{4}$ respectively. However, the RBF kernel shows higher SCs for $3\leq K\leq 8$. This seemingly irregular behaviour can be explained taking into consideration that the SC might be affected by the presence of macro and micro structures within the data, advantaging clustering with lower $K$ [40]. The SC of ZZ full also show an improvement as $K$ increases for all $\beta$ values, although significantly lower than ZZ linear case, suggesting that a more granular stratification better suits the data structure. The Z feature map, on the other hand, favours clustering with lower $K$ for $\beta\leq\frac{\pi}{2}$ being $K=5$ the best SC value. The trend then switches for $\beta>\frac{\pi}{2}$ for which higher values of $K$ yield a better SC. In addition, the experimental results demonstrate that the quality of the clustering, entailed by the SC, depends on the complexity of the encoding, i.e. on the entanglement structure of the feature map, and further by the angle $\beta$, which provides a regularisation of the expressivity. In fact, the entangled $ZZ$ feature maps strongly benefit from regularisation, as they show a subsequent decrease of the score with the increase of the angle range $\beta$. Whereas the Z feature map provides its bests results with some level of angle tuning for $\frac{\pi}{4}\leq\beta\leq\pi$ and has its pick for $\beta=\frac{\pi}{2}$.
@@ IV.1.1 Sample complexity analysis
To further assess the sample complexity of quantum models for our data, we focused on the best performing ZZ linear feature map. We selected the highest scoring embedding configuations by setting a threshold of $SC\geq 0.3$, and inspected how samples are grouped among clusters when the sample size varies from $1980$, the full set, to a subset of $N$ samples with $N\in\{50,100,250,500,800,1000\}$. We estimated clustering coherence of the reduced sets by calculating the AMI with respect to the full data for both classical RBF and quantum feature maps. The results are reported in Fig. 3. On average, the quantum derived clustering reach a plateau with an AMI $\geq 0.8$ at $N=250$ (Fig. 3 (a)) , slightly before the classical one. For the trivial case $K=2$ , which presents the highest SCs for both the classical and the quantum case, the $ZZ$ linear map with $\beta=\frac{\pi}{8}$ reaches perfect agreement (AMI = 1) at $N=100$ whereas the RBF kernel needs $N=500$ samples (Fig. 3 (b)). Similar outcomes can be observed with a more granular grouping of the samples with still high SC (i.e. $K=10$), in which the best configurations of quantum clustering achieve higher AMI compared to the classical one for each sample subset with $N<1000$. Moreover, the $ZZ$ linear with $\beta=\frac{\pi}{4}$, with the highest $SC=0.39$, shows a high coherence level with $AMI=0.9$ even with very few samples (i.e. $N=100$) (Fig. 3 (c)). This behaviour suggest that, for those quantum embeddings that achieve higher scores, the captured structure can be learnt with fewer data points.
@@ IV.2 Execution on QPU
Moving forward from the ideal simulation, we computed all QK on a real noisy NISQ device, to assess the reliability of numerical simulation in light of quantum noise, and verify how clusters might vary between different embeddings. The current limitations on available systems hinder the ability to run the experiments on the whole dataset, thus we consider, in the following, a reduced sampling of the data with $100$ points. This fact, while is expected to affect the results of the clustering analysis, does not undermine the capacity of addressing the overall effects of the quantum noise on the computational pipeline, as it depends mostly on the hardware specification and on the circuit composition.
@@
Feature map($U$) Entanglement type Circuit depth CX gate
@@
For each feature map we plotted the AMI results for each $K$ and $\beta$ values in Fig. 6. The deepest feature map (circuit depth $=63$), the ZZ linear, presents the lowest AMI scores, while the overall best AMI scores distribution are associated with the Z feature map (circuit depth $=5$), which is in principle the less expressive. The fully entangled feature map (circuit depth $=45$) shows intermediate values. Thus, the overall trend of the AMI confirms the previous findings in Fig 5: both the circuits depth and the gate composition emerging from the transpilation process, have a significant impact on the clusters consistency. In addition, using the Z feature map the coherence between simulated and real clusters increases with higher values of $\beta$: in particular, for $\beta>\frac{\pi}{4}$ and $K\leq 7$ , it achieves perfect agreement or almost perfect agreement. In contrast, the entangled feature maps exhibit the opposite trend, where a smaller $\beta$ corresponds to a higher AMI. By crossing these results from real quantum hardware with the SC obtained in the ideal scenario (Fig. 6 ), which provides the best possible outcome for these data, it is possible to draw some conclusion about the effect of quantum noise on the clustering. It can be noticed that quantum kernels of entangled feature maps with high $\beta$ give small SCs, thus they do not provide accurate clusters. In these cases the data can be easily partitioned very differently upon perturbations, potentially coming from any noisy source. Hence, when a real quantum hardware is in use, as there is an intrinsic noise source in the computation, the robustness of the clustering provided by the entangled kernel functions, with corresponding low SC values, is strongly hindered by the quantum noise.
V Conclusion
@@
We have investigated QK models for BC subtyping with a real multi-omics dataset addressing the challenges of problem-agnostic quantum embeddings for the best expressivity task-related trade-off. We explored three IPQ-style encoding schemes with different levels of entanglement of the qubits, comparing them with a classical RBF model. The QK obtained with a noiseless simulation on the full dataset ($N=1980$) show that the ZZ linear feature map achieved better clustering results, in terms of SC, compared to the classical RBF kernel for finer stratification of BC samples, with $K\geq 9$. In addition, varying the entanglement level and angle range, we noted how entangled feature map, which are in principle more expressive, strongly benefits from angle tuning, confirming that embedding optimisation is needed. Further, we focused on the best data-fitting feature map, the ZZ linear, and demonstrated experimentally that configurations with better clustering results ($SC\geq 0.3$) foster a reduction of sample complexity. Indeed for the best performing cases ($K=2,K=10$) the ZZ linear was able to obtain highly coherent stratification ($AMI\geq 0.9$) with the one obtained for the full dataset using only 100 samples. This finding, already identified by [27] in the supervised setting, is potentially highly valuable for clinical applications, where datasets and sample cohorts are often not sufficiently large and representative due to experimental and availability limitations. Finally, the execution on a NISQ system, the ibmq_Mumbai, showed that quantum noise has little to no effect on less expressive configurations, like the Z feature map that entails shallower circuits, while it hinders the results of entangled feature maps. This behaviour suggests that a trade-off between ideal performances and hardware resilience should be considered in quantum kernel selection. Nevertheless is important to note that NISQ devices now available are rapidly evolving and have reached higher accuracy level, compared to the QPU used in this work [44]. In addition, a major effort has been made also in advancing noise mitigation strategies[45], thus we expect that, in the foreseeable future, quantum protocols, like the one set out in this work, will achieve sufficient reliability and scalability to be implemented for real applications.
@@
The clustering of tumour samples using multi-omics data is crucial to identify clinically relevant subgroups, still the reconstruction of non linear relationship between tumoral molecular descriptors is made hard by classical model expressive capacity and limited data. Although QML and QK could offer an alternative view, we faced some major challenges in building proper quantum representation for biomedical systems, where formulating a problem-inspired embedding (eg. [46, 47, 48] ) is not possible. Nevertheless, similar considerations apply to classical ML, and given the early stage of quantum approaches, the heuristic exploration of classically complex problems is necessary to identify tasks that could be tackled with current NISQ devices. In this spirit, the present study aimed to investigate the potential use of NISQ devices and quantum algorithms in modeling multi-omics data for the specific task of tumour stratification.

### v2

@@ II.2 Quantum Embedding of classical data
[…]
@@ II.3 Expressivity regularization
The inherent high-dimensionality of the Hilbert space used for projecting the input data can hinder the model performance. In particular for QK it has been shown that, as a result of highly expressive problem agnostic encoding, the non-diagonal entries distribution often skews around the mean [35], making the points in the Quantum Hilbert space nearly indistinguishable. This phenomenon has a close parallelism to the curse of dimensionality of classical ML. It has been shown [36, 37] that acting on the parametric rotational gate angles can regularise the model expressivity. This strategy is implemented with a tunable hyperparameter called bandwidth, which is a constant within the range [0,1][0,1][ 0 , 1 ]. This parameter effectively scales the input, confining the embeddings to a more restricted region of the quantum state space as in:
@@ III.1 Breast Cancer Multi-omics dataset
III.2 Data encoding: quantum and classical
@@
The choice of a suitable quantum embedding, in case of complex real data, is not a straightforward task [11]. Nevertheless, it is possible to explore, as set out in this work, a sub set of configurations of the enconding scheme, in particular by varying the entanglement map of the unitary ansatz. In this way, a trade-off between the expressivity of the kernel and the task-related performance, can be assessed. We specifically explored three configuration of the IPQ-style encoding scheme, making use of unitary templates with different levels of entanglement of the qubits. The most simple configuration employed is the Z feature map without any entanglement, utilising a sequence of Hadamard and rotational $P$ gates, a single qubit rotation around the $Z$ axis, implemented in our case for a 4-dim classical input $x$ as shown in Fig. 1(a). We then implemented the ZZ feature map with both linear and full entanglement: in the linear entanglement, each qubit is entangled with its neighbouring qubits, while in the full entanglement, each qubit is entangled with every other qubit. These feature map resulted respectively in the embedding shown in Fig.1(b) and Fig. 1(c). Furthermore, we regularised the expressivity of the quantum embeddings tuning the angle range $\beta$. For each feature map, we tested five different scaling of the angle range so that the input data $x\in[0,\beta]$, where $\beta\in\{\frac{\pi}{8},\frac{\pi}{4},\frac{\pi}{2},\pi,2\pi\}$. To compare the quantum embeddings with a classical kernel we compute the RBF kernel, expressed by eq. 5, using the sklearn implementation with default parameter $\sigma=\sqrt{\frac{n}{2}}$ where $n$ is the number of feature [40]. We selected the RBF kernel due to its superior expressiveness compared to other classical kernels, enabling it to effectively capture complex, nonlinear relationships in the data [41] .
@@ III.3 Clustering evaluation
In which $a(i)$ and $b(i)$ are respectively the average distance of the $i$-th point to its own cluster, and the average distance to its second closest cluster. The value of the both the SC and the $s(i)$ ranges in the interval {$,1\}$ 1 } where a high value indicates that the object is well matched to its own cluster and weakly matched to neighbouring clusters. In addition, we employed Adjusted Mutual Information (AMI) as a measure of agreement between two sets of labelled data [45]. Given two partitions $U$ and $V$ of the same set of $N$ points, the AMI between them is calculated as:
@@ IV.1 Exact simulation
Figure 2: Geometric difference $g_{C|Q}$ between classical and quantum kernels of the full dataset (N=1980) plotted against the angle range $\beta$ and colored by feature map. The dashed gray line denotes the threshold for a significant difference between the kernels.
@@
[…]
@@ IV.1.1 Sample complexity analysis
To further assess the sample complexity of quantum models for our data, we focused on the best performing ZZ linear feature map. We selected the highest scoring embedding configuations by setting a threshold of $SC\geq 0.3$, and inspected how samples are grouped among clusters when the sample size varies from 198019801980, the full set, to a subset of $N$ samples with $N\in\{50,100,250,500,800,1000\}$. We estimated clustering coherence of the reduced sets by calculating the AMI with respect to the full data for both classical RBF and quantum feature maps. The results are reported in Fig. 4. On average, the quantum derived clustering reach a plateau with an AMI 0.8$\geq$ 0.8≥ 0.8 at $N=250$ (Fig. 4 (a)) , slightly before the classical one. For the trivial case $K=2$ , which presents the highest SCs for both the classical and the quantum case, the $ZZ$ linear map with $\beta=\frac{\pi}{8}$ reaches perfect agreement (AMI = 1) at $N=100$ whereas the RBF kernel needs $N=500$ samples (Fig. 4 (b)). Similar outcomes can be observed with a more granular grouping of the samples with still high SC (i.e. $K=10$), in which the best configurations of quantum clustering achieve higher AMI compared to the classical one for each sample subset with $N<1000$. Moreover, the $ZZ$ linear with $\beta=\frac{\pi}{4}$, with the highest $SC=0.39$, shows a high coherence level with $AMI=0.9$ even with very few samples (i.e. $N=100$) (Fig. 4 (c)). This behaviour suggest that, for those quantum embeddings that achieve higher scores, the captured structure can be learnt with fewer data points.
@@ IV.2 Execution on QPU
Moving forward from the ideal simulation, we computed all QK on a real noisy NISQ device, to assess the reliability of numerical simulation in light of quantum noise, and verify how clusters might vary between different embeddings. The current limitations on available systems hinder the ability to run the experiments on the whole dataset, thus we consider, in the following, a reduced sampling of the data with 100100100 points. This fact, while is expected to affect the results of the clustering analysis, does not undermine the capacity of addressing the overall effects of the quantum noise on the computational pipeline, as it depends mostly on the hardware specification and on the circuit composition.
@@
[…]
@@
For each feature map we plotted the AMI results for each $K$ and $\beta$ values in Fig. 7. The deepest feature map (circuit depth =63=63= 63), the ZZ linear, presents the lowest AMI scores, while the overall best AMI scores distribution are associated with the Z feature map (circuit depth =5=5= 5), which is in principle the less expressive. The fully entangled feature map (circuit depth =45=45= 45) shows intermediate values. Thus, the overall trend of the AMI confirms the previous findings in Fig 6: both the circuits depth and the gate composition emerging from the transpilation process, have a significant impact on the clusters consistency. In addition, using the Z feature map the coherence between simulated and real clusters increases with higher values of $\beta$: in particular, for $\beta>\frac{\pi}{4}$ and $K\leq 7$ , it achieves perfect agreement or almost perfect agreement. In contrast, the entangled feature maps exhibit the opposite trend, where a smaller $\beta$ corresponds to a higher AMI. By crossing these results from real quantum hardware with the SC obtained in the ideal scenario (Fig. 7 ), which provides the best possible outcome for these data, it is possible to draw some conclusion about the effect of quantum noise on the clustering. It can be noticed that quantum kernels of entangled feature maps with high $\beta$ give small SCs, thus they do not provide accurate clusters. In these cases the data can be easily partitioned very differently upon perturbations, potentially coming from any noisy source. Hence, when a real quantum hardware is in use, as there is an intrinsic noise source in the computation, the robustness of the clustering provided by the entangled kernel functions, with corresponding low SC values, is strongly hindered by the quantum noise.
IV.3 Limitations
Our results show that QKs achieve clustering performance comparable to the classical RBF kernel and a significant reduction in sample complexity, still these findings should be considered alongside certain limitations of the designed experimental set-up. Specifically, the outcomes are influenced by the data preprocessing step used to reduce dimensionality but its impact has not been investigated. Future studies should address this in terms of both reduction method and dimension of the projected space. In addition we acknowledge that, while METABRIC has been chosen as a benchmark for its multi-omics features, high dimensionality and heterogeneity, our results are dataset dependent and might not uphold to a different multi-omics tumor dataset. In particular, the relationship between the dataset characteristics and the observed advantages of the quantum model, like the reduction in sample complexity, was not quantitatively investigated as it goes beyond the scope of our work, but has emerged as an interesting aspect to study in the future in light of the practical application of QKs.
V Conclusion and Future work
@@
We have investigated QK models for BC subtyping with a real multi-omics dataset addressing the challenges of problem-agnostic quantum embeddings for the best expressivity task-related trade-off. We explored three IPQ-style encoding schemes with different levels of entanglement of the qubits, comparing them with a classical RBF model. The implemented QKs yield a $g_{C|Q}$ that consistently exceeds the threshold, thus guarantying a sufficient separation between the quantum and the classical models. The QK obtained with a noiseless simulation on the full dataset ($N=1980$) show that the ZZ linear feature map achieved better clustering results, in terms of SC, compared to the classical RBF kernel for finer stratification of BC samples, with $K\geq 9$. In addition, varying the entanglement level and angle range, we noted how entangled feature map, which are in principle more expressive, strongly benefits from angle tuning, confirming that embedding optimisation is needed. Further, we focused on the best data-fitting feature map, the ZZ linear, and demonstrated experimentally that configurations with better clustering results ($SC\geq 0.3$) foster a reduction of sample complexity. Indeed for the best performing cases ($K=2,K=10$) the ZZ linear was able to obtain highly coherent stratification ($AMI\geq 0.9$) with the one obtained for the full dataset using only 100 samples. This finding, already identified by [27] in the supervised setting, is potentially highly valuable for clinical applications, where datasets and sample cohorts are often not sufficiently large and representative due to experimental and availability limitations. While the proposed methodology showed promising results in crucial aspects for possible future applications of quantum models in biomedicine, we have tested here only a limited subset of possible encoding configurations. To further explore the encoding space to enhance performances, our work could be extended by using a trainable kernel approach [47]. In such case a parametrized QK is iteratively adapted to a dataset. In addition, we acknowledged that, our results are influenced by the preprocessing choices made for our data. Future work could investigate the impact of dimensionality reduction on performance, exploring both different reduction strategies and varying the dimensionality of the projected space (e.g. extending beyond the 4-dimensional space used in this study). Finally, the execution on a NISQ system, the ibmq_Mumbai, showed that quantum noise has little to no effect on less expressive configurations, like the Z feature map that entails shallower circuits, while it hinders the results of entangled feature maps. This behaviour suggests that a trade-off between ideal performances and hardware resilience should be considered in quantum kernel selection. Nevertheless is important to note that NISQ devices now available are rapidly evolving and have reached higher accuracy level, compared to the QPU used in this work [48]. In addition, a major effort has been made also in advancing noise mitigation strategies[49], thus we expect that, in the foreseeable future, quantum protocols, like the one set out in this work, will achieve sufficient reliability and scalability to be implemented for real applications.
@@
The clustering of tumour samples using multi-omics data is crucial to identify clinically relevant subgroups, still the reconstruction of non linear relationship between tumoral molecular descriptors is made hard by classical model expressive capacity and limited data. Although QML and QK could offer an alternative view, we faced some major challenges in building proper quantum representation for biomedical systems, where formulating a problem-inspired embedding (eg. [50, 51, 52] ) is not possible. Nevertheless, similar considerations apply to classical ML, and given the early stage of quantum approaches, the heuristic exploration of classically complex problems is necessary to identify tasks that could be tackled with current NISQ devices. In this spirit, we would like to further explore the use of QKs for challenging biomedical problems that could benefit from a quantum modeling of the system, such as peptide binding prediction, for which an embedding could be more easily formulated based on a-priori knowledge of the system.